A certain drug has a half-life in the body of 6.0h. What should the interval between doses be, if the concentration of drug in the body should not fall below

30.% of its initial concentration? Round your answer to 2 significant digits.
Answer #1

first order kinetics,

k = rate constant = 0.693/t1/2

t1/2 = half life = 6.0 h

k = 0.693/6 = 0.1155 hr-1

k = (1/t)ln(a0/a)

a0 = initial concentration of drug = 100%

a = remaining concentration of A after time t,hr = 30%

t = time = ? hr

0.1155 = (1/t)ln(100/30)

t = 10.42 hr

answer: Time interval = 1.0*10^1 hr